### v1

# Pseudo-Physics-Informed Neural Operators: Enhancing Operator Learning from Limited Data

Abstract
Neural operators have shown great potential in surrogate modeling. However, training a well-performing neural operator typically requires a substantial amount of data, which can pose a major challenge in complex applications. In such scenarios, detailed physical knowledge can be unavailable or difficult to obtain, and collecting extensive data is often prohibitively expensive. To mitigate this challenge, we propose the Pseudo Physics-Informed Neural Operator (PPI-NO) framework. PPI-NO constructs a surrogate physics system for the target system using partial differential equations (PDEs) derived from simple, rudimentary physics principles, such as basic differential operators. This surrogate system is coupled with a neural operator model, using an alternating update and learning process to iteratively enhance the model’s predictive power. While the physics derived via PPI-NO may not mirror the ground-truth underlying physical laws — hence the term “pseudo physics” — this approach significantly improves the accuracy of standard operator learning models in data-scarce scenarios, which is evidenced by extensive evaluations across five benchmark tasks and a fatigue modeling application.
Machine Learning, ICML
1 Introduction
Operator learning, an important area for data-driven surrogate modeling, has made significant strides with the emergence of neural operators, which leverage the expressive power of neural networks. Notable examples include Fourier Neural Operators (FNO) (Li et al., 2020b, ), Deep Operator Net (DONet) (Lu et al.,, 2021) and other frameworks such as (Cao,, 2021; Hao et al.,, 2023). FNO employs Fourier transform for global convolution and function transformation, while DONet introduces two sub-networks, the branch net and trunk net, to extract representations from the functional space and query locations, respectively, enabling predictions akin to attention mechanisms (Vaswani et al.,, 2017).
For trading for model capacity and performance, neural operators typically require a substantial amount of training data to perform optimally. This demand poses challenges, particularly in complex problems, where training data is limited and costly to acquire. In response, the field of physics-informed machine learning, including physics-informed neural networks (PINN) (Raissi et al.,, 2019), has shown promise by incorporating physical laws as soft constraints during training. This approach serves as a regularization technique, embedding a fundamental understanding of physics into the model to lessen its reliance on extensive data. Building on this idea, the concept of physics-informed neural operators (PINO) has emerged, which integrates physical laws as soft constraints to enhance operator learning while reducing data quantity. It has been used in (Wang et al.,, 2021; Li et al.,, 2021) for FNO and DONet training.
Despite the success of PINO, the necessity for a thorough understanding of the underlying physics can pose a significant hurdle, especially in complex applications such as in fracture mechanics and climate modeling. In those scenarios, the detailed physical knowledge is often unavailable or difficult to identify, and it is often prohibitively expensive to collect extensive data. To navigate these challenges while retaining the benefits of physics-informed learning, we propose the Pseudo Physics-Informed Neural Operator (PPI-NO). This framework bypasses the need for exhaustive physical comprehension by constructing a neural-network-based partial differential equation (PDE) that characterizes the target system directly from data. The neural PDE is then coupled with the neural operator for alternating updates and training, enabling iterative extraction, refinement and integration of physics knowledge to enhance operator learning. The contribution mainly lies in the following aspects:
  • To our knowledge, PPI-NO is the first work to enhance a standard operator learning pipeline using physics directly learned from limited data, delivering superior accuracy without the need for in-depth physical understanding or extensive data collection.
  • PPI-NO opens up a new paradigm of physics-informed machine learning where only rudimentary physics assumptions (in this case, the basic differential operations) are required rather than in-depth or rigorous expert knowledge, extending the spectrum of the physics-informed learning for experts of different levels.
  • The effectiveness of PPI-NO is validated through extensive evaluations on five commonly used benchmark operator learning tasks in literature (Li et al., 2020b, ; Lu et al.,, 2022), including Darcy flow, nonlinear diffusion, Eikonal, Poisson and advection equations, as well as one application in fatigue modeling in fracture mechanics, where the ground-truth holistic PDE system is unknown.
2 Background
Problem Formulation. Operator learning seeks to approximate an operator that maps input parameters and/or functions to corresponding output functions. In many cases, operator learning rises in the context of solving partial differential equations (PDEs), where the operator corresponds to the solution operator of the PDE. Consider a PDE system:
$\mathcal{N}[u](\mathbf{x})=f(\mathbf{x}),\quad\mathbf{x}\in\Omega\times[0,\infty),$ (1)
where ${\bf x}$ is a compact notation for the spatial and temporal coordinates, $\Omega$ is the spatial domain, $[0,\infty)$ is the temporal domain, $\mathcal{N}$ is a nonlinear differential operator, $u(\mathbf{x})$ is the solution function, and $f(\mathbf{x})$ is the source term. We aim to learn the solution operator of the PDE system, ${\psi}:\mathbb{F}\rightarrow\mathbb{U}$ where $\mathbb{F}$ and $\mathbb{U}$ are two functional spaces, using a training dataset $\mathcal{D}=\{({\bf f}_{n},{\bf u}_{n})\}_{n=1}^{N}$, which includes different instances of $u(\cdot)$ and $f(\cdot)$ sampled/discretized at a set of locations. Once the model is trained, it can be used to directly predict the solution function $u$ for new instances of the input $f$, offering a much more efficient alternative to running numerical solvers from scratch. However, the training dataset still needs to be generated offline using numerical solvers.
[…]
Deep Operator Network (DONet) (Lu et al.,, 2021) is another prominent work in operator learning. The architecture of a DONet is structured into two components: a branch net and a trunk net, learning representations for the input functions and querying locations, respectively. Consider an input function $f({\bf x})\in\mathbb{F}$ sampled at $m$ locations $\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{m}\}$ and an output function $u\in\mathbb{U}$. The branch net receives the values $[f({\bf x}_{1}),f({\bf x}_{2}),\cdots,f({\bf x}_{m})]$ and outputs a feature representation $[b_{1},b_{2},\cdots,b_{p}]^{\top}\in\mathbb{R}^{p}$. Concurrently, the trunk network processes a querying location ${\bf x}$ and outputs another feature vector $[t_{1},t_{2},\cdots,t_{p}]^{\top}\in\mathbb{R}^{p}$. The output function value at ${\bf x}$ is predicted as a sum of products of the corresponding elements from the branch and trunk nets, $\mathcal{G}[f]({\bf x})\approx\sum_{k=1}^{p}b_{k}t_{k}$, where $\mathcal{G}$ is the learned operator mapping $f$ to $u$.
Physics-Informed Neural Operator (PINO) (Wang et al.,, 2021; Li et al.,, 2021) has recently emerged as a promising approach to address the data scarcity issue in operator learning. PINO embeds physical laws — typically governing equations — into the learning process. The incorporation of physics not only enhances the model’s adherence to ground-truth phenomena but also reduces its dependency on extensive training data. Mathematically, the integration of physics into the learning process can be viewed as adding a regularization term in the loss function. Let $\mathcal{L}_{\text{data}}$ represent the standard data-fitting loss term (e.g., the mean squared error between the predicted and actual outputs), and the physics-informed term $\mathcal{L}_{\text{physics}}$ be the residual of the governing PDEs evaluated at the neural network’s outputs. The loss function $\mathcal{L}$ for a PINO model is expressed as
$\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}},$
where $\lambda$ is a weighting factor that balances the importance of data-fitting versus physics compliance. This approach encourages the model to learn solutions not only consistent with the provided data but also physically plausible.
3 Methodology
In the absence of physics knowledge (i.e., PDE system (1)), it is impossible to construct the physics loss term as in PINO. To address this challenge, we propose a “pseudo” physics-informed operator learning framework that extracts useful physics representation from data so as to enhance operator learning. This framework is motivated by relatively complex applications, where data is often costly and/or limited while the underlying physics is hard to fully understand.
3.1 Pseudo Physics System Learning
As the first step, we propose a novel approach to learn the physics system using scarce training data. Our key observation is that, although the mapping from $f$ to $u$ can be intricate and may necessitate information across the entire domain (e.g., in linear PDEs, $u$ is an integration of the Green’s function multiplied with $f$ over the domain), the underlying PDE system (1) simplifies to a local combination of $u$ and its derivatives. We therefore design a neural network $\phi$ to approximate the general form of $\mathcal{N}$,
$\mathcal{N}[u]({\bf x})\approx\phi\left({\bf x},u({\bf x}),S_{1}(u)({\bf x}),\ldots,S_{Q}(u)({\bf x})\right),$ (2)
where $\{S_{j}\}_{j=1}^{Q}$ are $Q$ derivative operators that we believe should be present in the system, such as $\partial_{t}u$, $\partial_{tt}u$, $\partial_{x_{1}}u$, $\partial_{x_{2}}u$, $\partial_{x_{1}x_{1}}u$, $\partial_{x_{1}x_{2}}u$, $\partial_{x_{2}x_{2}}u$, and more.
The inherent local combinatorial nature of the PDE representation decouples the values of $u$ and its derivatives across different sampling locations, thereby significantly reducing the learning complexity and the amount of training data required. For instance, consider sampling the input function $f$ and output function $u$ on a $128\times 128$ grid. A single pair of discretized input and output functions, denoted as $({\bf f},{\bf u})$, is typically insufficient for a neural operator to effectively learn the mapping $f\rightarrow u$. However, this pair can be decomposed into $128\times 128=16,384$ training data points across various (spatial and temporal) locations, which can be used to train $\phi$ as outlined in (3). This decomposition provides rich information that reveals the local relationships between those derivatives. Therefore, even with a small number of $({\bf f},{\bf u})$ pairs, we hypothesize that the learning of the PDE system $\mathcal{N}$ through our formulation in (3) can still yield high accuracy in predicting $f(\cdot)$ as in (1).
Figure 1: The illustration of “pseudo” physics representation network $\phi$. The input consists of $u$ and its finite difference derivative approximations ($\{\hat{S}_{1}(u),\ldots,\hat{S}_{Q}(u)\}$) across different sampling locations. The top row shows a convolution layer that aggregates local neighborhood to compensate the information loss caused by finite difference. The bottom row shows that $\phi$ uses fully connected layers at each sampling location to combine $u$ and its derivatives locally to predict $f$ at the same location.
Figure 2: PPI-NO learning framework.
We use an $L_{2}$ loss to estimate the parameters of $\phi$, which is defined as
${\mathcal{L}}_{\phi}=\sum_{n=1}^{N}\sum_{j=1}^{M}$ $\Big{[}\phi({\bf x}_{j},u_{n}({\bf x}_{j}),S_{1}(u_{n})({\bf x}_{j}),\ldots,S_{Q}(u_{n})({\bf x}_{j}))$
$-f_{n}({\bf x}_{j})\Big{]}^{2},$ (3)
where $f_{n}(\cdot)$ and $u_{n}(\cdot)$ are the input and output functions in $n$-th training example, and $\{{\bf x}_{1},\ldots{\bf x}_{M}\}$ are the locations at which we sample $f_{n}$ and $u_{n}$.
We use finite difference to obtain the derivatives of each $u_{n}$, namely, $S_{k}(u_{n})$ ($1\leq k\leq Q$), and then pass these inputs to the network $\phi$ to compute the prediction. Since finite difference introduces numerical errors, the computed derivatives are not exact. To compensate for such information loss, we propose incorporating a convolution layer to integrate and leverage the neighborhood information (see Fig. 1 top). Specifically, let $\hat{S}_{k}(u_{n})$ represents the finite difference approximation of $S_{k}(u_{n})$ ($1\leq k\leq Q)$; for $k=0$, we define $\hat{S}_{k}(u_{n})=u_{n}$. Each $\hat{S}_{k}(u_{n})$ is treated as an input channel. After applying the convolution, the output at each sampling location ${\bf x}_{j}$ for channel $k$ is given by
$\overline{S}_{k}({\bf x}_{j})=\sum_{c=0}^{Q}\sum_{{\bf x}_{i}\in\text{nei}({\bf x}_{j})}w_{c}(\Delta_{ij})\hat{S}_{c}({\bf x}_{j}),$ (4)
where $\text{nei}(\cdot)$ denotes the neighborhood sampling locations defined in the convolution filter, $\Delta_{ij}$ is the relative distance between ${\bf x}_{i}$ and ${\bf x}_{j}$, and $w_{c}(\Delta_{ij})$ is the corresponding filter weight. We view $\overline{S}_{k}({\bf x}_{j})$ as an interpolation, providing a new approximation of $S_{k}({\bf x}_{j})$ by incorporating all neighborhood values of $u_{n}({\bf x}_{j})$ and their finite difference derivative approximations. The convolution results are then passed into subsequent layers. The interpolation (filter) weights are jointly learned from data. This design allows us to integrate additional information into the inputs to facilitate the learning of $\phi$. Empirical results confirm that incorporating this convolution layer improves the prediction accuracy; refer to the ablation study in Section 5.1 and Appendix Table 6(a). Next, at each sampling location ${\bf x}_{j}$, we employ fully connected layers (i.e., MLP) to combine all $\{\overline{S}_{k}({\bf x}_{j})\}_{k=0}^{Q}$ in an arbitrarily flexible way to predict $f({\bf x}_{j})$; see Fig. 1 bottom. In this way, the learned neural network mapping $\phi:u\rightarrow f$, although black-box in nature, is expressive enough to encapsulate valuable physics knowledge inherent in the data employed for operator learning.
Our method can be readily adapted to other numerical approaches for derivative approximation. For instance, when functions are irregularly sampled, smooth function estimators such as kernel interpolation (Long et al.,, 2024), the RBF-FD method (Fornberg et al.,, 2013; Fornberg and Flyer,, 2015), or Bayesian B-splines (Sun et al.,, 2022) can be used to estimate gradient information directly from data. These estimated gradients are then provided as inputs to our PDE neural network $\phi$ for further training.
3.2 Coupling Neural Operator with Pseudo Physics
Next, we leverage the pseudo physics laws embedded in the learned mapping $\phi:u\rightarrow f$ to enhance neural operator learning. Specifically, we use $\phi$ to reconstruct $f$ from the $u$ predicted by the neural operator. In this way, our approach goes beyond relying solely on the training data; it uses the physics learned in the previous step to incorporate a reconstruction error into the learning of the neural operator parameters; see Fig. 2 for an illustration.
Initially, we train the neural operator $\psi:f\rightarrow u$ using the available training data, creating a preliminary model. This model is developed using FNO or DONet or other neural operators. The focus is to first establish a basic understanding of the relationship between $f$ and $u$ from the limited data. Next, the loss function for $\psi$ is augmented using the physics laws learned in the first step,
$\mathcal{L}=\sum_{n=1}^{N}\mathcal{L}_{2}(\psi(f_{n}),u_{n})+\lambda\cdot\mathbb{E}_{p(f^{\prime})}\left[\mathcal{L}_{2}(f^{\prime},\phi(\psi(f^{\prime})))\right],$
where the first term is the $\mathcal{L}_{2}$ loss for data fitting (as in the standard neural operator training), and the second term is the expected reconstruction error for the input function. The second term incorporates the physics laws embedded in $\phi(\cdot)$, and $\lambda$ is a weighting factor that balances the training data loss against the reconstruction error.
In practice, the expected reconstruction error does not have a closed form. One can sample a collection of $f^{\prime}$ from the underlying distribution of the input function $p(\cdot)$, e.g., a Gauss random field or Gaussian process, and then employ a Monte-Carlo approximation,
$\mathcal{L}=\sum_{n=1}^{N}\mathcal{L}_{2}(\psi(f_{n}),u_{n})+\lambda\frac{1}{N^{\prime}}\sum_{n=1}^{N^{\prime}}\mathcal{L}_{2}(f^{\prime}_{n},\phi(\psi(f^{\prime}_{n}))),$ (5)
where $N^{\prime}$ is the number of input function samples.
To enhance the operator learning process, the model is iteratively refined. In each iteration, we first fine-tune the neural operator $\psi$ with the pseudo physics $\phi$ fixed, and then fix $\psi$, fine-tune $\phi$ to refine the physics representation. This fine-tuning loop is carried out for multiple iterations, allowing for continuous improvement of the neural operator based on the refined physics representation. Our method is summarized in Algorithm 1.
Algorithm 1 Pseudo-Physics-Informed NO
1:  Train a preliminary NO $\psi$ with standard NO loss.
2:  Train a preliminary psuedo physics network $\phi$ with loss ${\mathcal{L}}_{\phi}$ in (3).
3:  repeat
4:     Sample $N^{\prime}$ instances from the input function space.
5:     Fix pseudo physics network $\phi$, fine tune NO $\psi$ with the loss (5).
6:     Fix NO $\psi$, fine tune the pseudo physics network $\phi$ with the loss (5).
7:  until Maximum iterations are done or convergence
This methodology mirrors human experts’ approach to physics system modeling, where a limited amount of data is available and used to investigate the physics laws inspired by simple differential operations, and then these laws are utilized to generalize the system for further applications, e.g., forward simulation and solving inverse problems. In our method, the reconstruction loss term augments the operator learning with additional information, leading to potential improvement upon only training with limited data.
4 Related Work
Operator learning is a rapidly evolving research field, with a variety of methods categorized as neural operators. Alongside FNO, several notable approaches have been introduced, such as low-rank neural operator (LNO) (Li et al., 2020b, ), multipole graph neural operator (MGNO) (Li et al., 2020a, ), multiwavelet-based NO (Gupta et al.,, 2021), and convolutional NOs (CNO) (Raonic et al.,, 2023). Deep Operator Net (DON) (Lu et al.,, 2021) is another popular approach, consisting of a branch network applied to input function values and a trunk network applied to sampling locations. The final prediction is obtained through the dot product of the outputs from the two networks. To enhance stability and efficiency, Lu et al., (2022) proposed replacing the trunk network with POD (PCA) bases. Recently, transformer architectures have been employed to design neural operators, offering a promising direction for operator learning, e.g., (Cao,, 2021; Li et al.,, 2022; Hao et al.,, 2023).
Physics-Informed Neural Networks (PINNs) (Raissi et al.,, 2019) mark a significant advancement in scientific machine learning. PINNs integrate physical laws directly into the learning process, making them effective for solving differential equations and understanding complex physical systems. This methodology is particularly beneficial in scenarios where data is limited or expensive to obtain. Pioneering the concept of PINO, Li et al., (2021) introduced a dual-resolution approach that combines low-resolution empirical data with high-resolution PDE constraints. This method achieves precise emulation of solution operators across various PDE classes. In parallel, physics-informed DONet by Wang et al., (2021) incorporate regularization strategies enforcing physical law adherence into the training of DONets. Zanardi et al., (2023) presented an approach using PINO for simulations in non-equilibrium reacting flows.  Lee et al., (2023) proposed opPINN, a framework combining physics-informed neural networks with operator learning for solving the Fokker-Planck-Landau (FPL) equation.  Rosofsky et al., (2023) provided a review of applications of physics-informed neural operators. However, existing methods demand one should know the physics laws beforehand, which might not be feasible in many practical applications or complex systems. Our method offers a simple and effective framework, enabling the extraction of implicit physics laws directly from data, even when the data is scarce. Empirically, these pseudo physics laws have proven to be highly beneficial in enhancing the performance of operator learning, as demonstrated in Section 5. Many methods have been developed specifically for discovering differential equations from data, including SINDy (Brunton et al.,, 2016) and its variants (Schaeffer,, 2017; Zhang and Ma,, 2020; Lagergren et al.,, 2020), PINN-SR (Chen et al.,, 2021), Bayesian spline learning (Sun et al.,, 2022), and kernel-based equation discovery (Long et al.,, 2024). To ensure interpretability, these approaches typically assume a specific equation form and perform sparse selection from a set of candidate operators. In contrast to these methods, which prioritize interpretability, our approach focuses on enhancing the prediction accuracy of operator learning under limited data. To this end, we utilize a black-box neural network to represent PDEs. While this offers greater flexibility, it comes at the cost of reduced interpretability. Our method employs an alternating update strategy to jointly refine the PDE representation and improve operator learning.
Our work is also related to the cycle consistence framework (Zhu et al.,, 2017) for image-to-image translation. A critical difference is that cycle-consistence performs unpaired image-to-image translation, while our method aims for accurate paired translation (mapping). In cycle-consistence, the translation is viewed successfully as long as the translated images follow the target distribution. Hence, cycle-consistence has a much more relaxed objective. Another key difference is that our method aims to improve the learning of a function-to-function mapping with very limited data— that motivates us to learn a “pseudo” physics representation. The cycle-consistence framework relies on adversarial training which typically requires a large amount of data to obtain successful learning outcomes.
5 Experiments
We tested on five commonly used benchmark operator learning problems in literature (Li et al., 2020b, ; Lu et al.,, 2022), including Darcy Flow, Nonlinear Diffusion, Eikonal, Poisson and Advection. In addition, we examined our method in an application of fatigue modeling. The task is to predict the stress intensity factor (SIF) for semi-elliptical surface cracks on plates, given three geometric parameters that characterize the cracks (Merrell et al.,, 2024); see Appendix Fig. 4. The SIF plays a critical role in modeling crack growth by quantifying the stress state near the tip of a crack, and hence SIF computation and analysis is extremely important in fatigue modeling and fracture mechanics (Anderson and Anderson,, 2005). The SIF computation is expensive, because it typically needs to run finite element method (FEM) or extended FEM with very fine meshes (Kuna,, 2013). Due to the complex sequence of computational steps involved in SIF calculation, there is no holistic PDE that directly models the relationship between the geometric features and the SIF function. Instead, SIF computation typically relies on numerical methods and the extraction of local stress fields near the crack tip. The details about all the datasets are given in Section A of the Appendix.
We evaluated our method based on two widely used NO models, FNO and DONet. Note that our method is straightforward to implement on other NO models, such as attention based models.
For each operator learning benchmark, we simulated 100 examples for testing, and varied the number of training examples from {5, 10, 20, 30}, except for Advection, we ran with {20, 30, 50, 80} training examples. For SIF prediction, which is much more challenging, we experimented with training size from {400, 500, 600}, and employed 200 test examples. We repeated the evaluation for five times, each time we randomly sampled a different training set. The input to the pseudo physics networks $\phi$ includes all the partial derivatives up to the 2nd order. The details about the hyperparameter settings are given in Appendix Section A.7.
Table 1: Relative $L_{2}$ error in five operator learning benchmarks, where “PPI” is short for “Pseudo-Physics Informed”. The results were averaged from five runs.
(a) Darcy flow
Training size 5 10 20 30
FNO 0.4915 $\pm$ 0.0210 0.3870 $\pm$ 0.0118 0.2783 $\pm$ 0.0212 0.1645 $\pm$ 0.0071
PPI-FNO 0.1716 $\pm$ 0.0048 0.0956 $\pm$ 0.0084 0.0680 $\pm$ 0.0031 0.0642 $\pm$ 0.0010
Error Reduction 65.08% 75.29% 75.56% 60.97%
DONet 0.8678 $\pm$ 0.0089 0.6854 $\pm$ 0.0363 0.5841 $\pm$ 0.0279 0.5672 $\pm$ 0.0172
PPI-DONet 0.5214 $\pm$ 0.0543 0.3408 $\pm$ 0.0209 0.2775 $\pm$ 0.0224 0.2611 $\pm$ 0.0084
Error Reduction 39.91% 50.27% 52.49% 53.96%
(b) Nonlinear diffusion
Training size 5 10 20 30
FNO 0.2004 $\pm$ 0.0083 0.1242 $\pm$ 0.0046 0.0876 $\pm$ 0.0061 0.0551 $\pm$ 0.0021
PPI-FNO 0.0105 $\pm$ 0.0016 0.0066 $\pm$ 0.00023 0.0049 $\pm$ 0.00037 0.0038 $\pm$ 0.00039
Error Reduction 94.76% 94.68% 94.40% 93.10%
DONet 0.3010 $\pm$ 0.0119 0.2505 $\pm$ 0.0057 0.1726 $\pm$ 0.0076 0.1430 $\pm$ 0.0036
PPI-DONet 0.1478$\pm$ 0.0126 0.1161 $\pm$ 0.0124 0.1032 $\pm$ 0.0059 0.0842 $\pm$ 0.0041
Error Reduction 50.89% 53.65% 40.20 % 41.11%
(c) Eikonal
Training size 5 10 20 30
FNO 0.2102 $\pm$ 0.0133 0.1562 $\pm$ 0.0098 0.0981 $\pm$ 0.0022 0.0843 $\pm$ 0.0020
PPI-FNO 0.0678 $\pm$ 0.0026 0.0582 $\pm$ 0.0043 0.0493 $\pm$ 0.0023 0.0459 $\pm$ 0.0010
Error Reduction 67.74% 62.74% 49.74% 45.55%
DONet 0.3374 $\pm$ 0.0944 0.1759 $\pm$ 0.0065 0.1191 $\pm$ 0.0047 0.1096 $\pm$ 0.0037
PPI-DONet 0.1302$\pm$ 0.0127 0.0907 $\pm$ 0.0093 0.0714 $\pm$ 0.0011 $0$$\pm$ 0.0007
Error Reduction 61.41% 48.43% 40.05% 36.13%
(d) Poisson
Training size 5 10 20 30
FNO 0.2340 $\pm$ 0.0083 0.1390 $\pm$ 0.0007 0.0895 $\pm$ 0.0008 0.0698 $\pm$ 0.0014
PPI-FNO 0.1437 $\pm$ 0.0062 0.0771 $\pm$ 0.0018 0.0544 $\pm$ 0.0009 0.0458 $\pm$ 0.0003
Error Reduction 38.59% 44.53% 39.22% 34.38%
DONet 0.6142 $\pm$ 0.0046 0.5839 $\pm$ 0.0090 0.5320 $\pm$ 0.0028 0.5195 $\pm$ 0.0040
PPI-DONet 0.5275 $\pm$ 0.0037 0.5001 $\pm$ 0.0042 0.4450 $\pm$ 0.0010 0.4258 $\pm$ 0.0040
Error Reduction 14.12% 14.35% 16.35% 18.04%
(e) Advection
Training size 20 30 50 80
FNO 0.4872 $\pm$ 0.0097 0.4035 $\pm$ 0.0086 0.3019 $\pm$ 0.0085 0.2482 $\pm$ 0.0059
PPI-FNO 0.3693 $\pm$ 0.0099 0.3224 $\pm$ 0.0123 0.2236 $\pm$ 0.0075 0.1698 $\pm$ 0.0075
Error Reduction 24.20% 20.10% 25.94% 31.59%
DONet 0.5795 $\pm$ 0.0045 0.4810 $\pm$ 0.0092 0.3882 $\pm$ 0.0086 0.3164 $\pm$ 0.0072
PPI-DONet 0.3630 $\pm$ 0.0112 0.2897 $\pm$ 0.0097 0.2629 $\pm$ 0.0053 0.2120 $\pm$ 0.0065
Error Reduction 37.36% 39.77% 32.28% 33.00%
5.1 Results and Analysis
(a) PPI-FNO: learning
(b) PPI-DONet: learning
(c) PPI-FNO: $\lambda$
(d) PPI-DONet: $\lambda$
Figure 3: Learning curve of PPI-FNO on Darcy Flow (a), and of PPI-DONet on nonlinear diffusion (b). In (c) and (d) we show how the weight $\lambda$ of “pseudo physics” affects the operator learning performance. The horizontal line in (c) and (d) are the relative $L_{2}$ errors of standard FNO and DONet.
Table 2: SIF prediction error for plate surface cracks in fatigue modeling.
Training size 400 500 600
FNO 0.1776 $\pm$ 0.0150 0.1695 $\pm$ 0.0090 0.1122 $\pm$ 0.0094
PPI-FNO 0.1166 $\pm$ 0.0064 0.1151 $\pm$ 0.0093 0.0850$\pm$0.0060
Error Reduction 34.35% 32.09% 24.24%
DONet 0.5318 $\pm$ 0.0095 0.5155 $\pm$ 0.0200 0.4037 $\pm$ 0.0331
PPI-DONet 0.3490 $\pm$ 0.0034 0.3468 $\pm$ 0.0074 0.3299 $\pm$ 0.0066
Error Reduction 34.37% 32.73% 18.28%
Predictive performance with limited data. We reported the average relative $L_{2}$ error and the standard deviation (with and without using pseudo physics informed learning) in Table 1 and Table 2. The model trained with the pseudo physics network (see (5)) is denoted as PPI-FNO or PPI-DONet, short for Pseudo Physics Informed FNO/DONet. In all the cases, with our pseudo physics informed approach, the prediction error of both FNO and DONet experiences a large reduction. For instance, across all training sizes in Darcy Flow and nonlinear diffusion, PPI-FNO reduces the relative $L_{2}$ error of the ordinary FNO by over 60% and 93%, respectively. In Darcy Flow with training sizes 10 to 30, PPI-DONet reduces the error of the ordinary DONet by over 50%. In SIF prediction, our method applied to both FNO and DONet reduced the error by over 30% with training size 400 and 500. Even the minimum reduction across all cases achieves 14.12% (PPI-DONet over DONet on Poisson with training size 5).
Together these results demonstrate the strong positive impact of the learned physics by our neural network model $\phi$ specified in Section 3.1. Although it remains opaque and non-interpretable, it encapsulates valuable knowledge that greatly enhances the performance of operator learning with limited data.
Next, we assessed the accuracy of the learned physics laws by examining the relative $L_{2}$ error in predicting the source functions $f$ from $\phi$ (see (3)). We tested on Darcy Flow, nonlinear diffusion, and Eikonal. We compared a baseline method that removes the convolution layer of $\phi$, leaving only the fully connected layers, namely MLP (see Fig 1 bottom). The results are reported in Appendix Table 5. It can be observed that in nearly every case, adding a convolution layer indeed significantly improves the accuracy of $\phi$. This improvement might be attributed to the convolution layer’s ability to integrate neighboring information and compensate for the information loss introduced by finite difference in approximating the derivatives. We also experimented with multiple convolution layers, but the improvement was found to be marginal.
In addition, we found the operator learning improvement is relatively robust to the accuracy of our physics representation $\phi$. For instance, on Darcy Flow with training size 5 and 10, the relative $L_{2}$ error of $\phi$ network is 0.2285 and 0.1392, which is significantly bigger than with training size 30 where the relative $L_{2}$ error is 0.0688. Yet the error reduction upon FNO (see Table 3(a)) under all the three training sizes is above 60%. The error reduction upon DONet is 40% for training size 5 and over 50% for training size 10 and 30. The results imply that even roughly capturing the underlying physics (with $\phi$) can substantially boost the operator learning performance. In Appendix Section B, we we provide a detailed comparison of the point-wise error of our method against standard NO, along with additional discussions.
Training with rich data. Although our method is designed to enhance operator learning with limited data, we also evaluated its performance when the training data is abundant. Specifically, we increased the training size of the five operator learning benchmarks to 600 and 1000 examples. The results, presented in Appendix Table 3, show that the accuracy of the standard NO and our method, PPI-NO, becomes comparable. For Darcy flow and Advection, our method achieves slight improvement, whereas for Eikonal and Nonlinear diffusion, it performs slightly worse. This may be because, with sufficiently rich data, the standard NO can already utilize the data effectively to achieve good performance, leaving little room for the added value of the learned physics to further improve the results. In such cases, the benefit of incorporating physics knowledge becomes marginal. Furthermore, our alternating updating mechanism brings additional optimization workload, which may introduce additional complexity to the learning process.
Comparison with NO training using ground-truth physics. Our work is motivated by scenarios where the underlying physics is unknown. It is therefore insightful to compare our approach with training NO using the ground-truth physics, as done in PINO. This comparison allows us to examine how our learned “pseudo” physics differs from the true physics in facilitating operator learning. To this end, we tested PINO on the Poisson and Advection cases, using finite difference approximations to compute the derivatives. We used FNO as the baseline. The PDE residual was incorporated into the training loss. As shown in Appendix Table 4, while PINO consistently achieves even better results, the performance of our method is close. This demonstrates that our learned black-box representation, despite lacking interpretability, can provide a similar benefit in improving operator learning performance.
Learning behavior. We examined the learning behavior of our method, which conducts an iterative, alternatingly fine-tuning process. We employed one Darcy Flow, one nonlinear diffusion and one Eikonal dataset, each with 30 examples. We show the test relative $L_{2}$ error along with the iterations in Fig. 3(a), 3(b), and Appendix Fig. 6(a) and Fig. 6(b). As we can see, the predictive performance of our algorithm kept improving and tended to converge at last, affirming the efficacy of the learning process.
Ablation study on pseudo physics network $\phi$. To confirm the efficacy of our designed network $\phi$ in facilitating operator learning, we considered alternative designs for $\phi$: (1) using standard FNO to predict $f$ directly from $u$; no derivative information is included in the input; (2) removing the convolution layer in our model, and just keeping the fully-connected layers, namely MLP; the derivative information is not included in the input. With different designs of $\phi$, we evaluated the PPI learning performance on the Darcy Flow benchmark. The relative $L_{2}$ error in predicting $f$ via $\phi$ and predicting $u$ is reported in Appendix Table 6. It can be seen that our design of $\phi$ consistently outperforms alternative architectures by a notable margin, showing the effectiveness of learning a (black-box) PDE representation and improving the operator learning.
Ablation study on the choice of derivatives. We further investigated the PPI learning performance with respect to the choice of derivatives used in our pseudo physics network. Specifically, we tested PPI-FNO on the Darcy-flow benchmark and varied the order of derivatives up to 0, 1, 2, and 3. The performance is reported in Appendix Table 7. We can see that although the accuracy of $\phi$ with derivatives up to the third order is slightly better than with derivatives up to the second order, the best operator learning performance was still achieved using derivatives up to the second order (which was used in our evaluations). This might be because higher-order derivative information can cause overfitting in the pseudo physics network $\phi$ to a certain degree. Such higher-order information may not be critical to the actual mechanism of the physical system and can therefore impede the improvement of operator learning performance.
Ablation study on the weight $\lambda$. We examined the effect of the weight $\lambda$ of our “pseudo physics”; see (5). To this end, we used one Darcy Flow dataset and nonlinear diffusion dataset with training size 30. We varied $\lambda$ from $[0.5,10^{2}]$, and run PPI-FNO and PPI-DONet on Darcy Flow and nonlinear diffusion, respectively. As shown in Fig. 3(c) and 3(d), we can see that across a wide range of $\lambda$ values, PPI-FNO and PPI-DONet can consistently outperform the standard FNO and DONet respectively by a large margin. However, the choice $\lambda$ does have a significant influence on the operator learning performance, and the best choice is often in between. In Appendix Fig. 6(c) and 6(d), we show results on Ekonal, which we make similar observations.
Memory Cost. Our PPI-NO is more costly than standard NO since we need to train an additional “pseudo” physics network $\phi$ along with the NO model. However, the network $\phi$ is small as compared to the NO component — $\phi$ is simply a pixel-wise MLP coupled with one convolution filter, resulting in a marginal increase in memory cost. Appendix Table 8 shows the parameter count of FNO, DONet and their pseudo-physics-informed version. On average, PPI-FNO increases the number of parameters over FNO by 1.29% while PPI-DONet over DONet by 1.89%.
6 Conclusion
We have presented a Pseudo Physics-Informed Neural Operator (PPI-NO) learning framework. PPI-NO is based on our observation that a PDE system is often characterized by a local combination of the solution and its derivatives. This property makes it feasible to learn an effective representation of the PDE system, even with limited data. While the physics delineated by PPI-NO might not precisely reflect true physical phenomena, our findings reveal that this method significantly enhances the efficiency of operator learning with limited data quantity.
7 Impact Statements
This paper presents work whose goal is to advance machine learning for operator learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix
Appendix A Experimental Details
A.1 Darcy Flow
We considered a steady-state 2D Darcy Flow equation (Li et al., 2020b, ),
$-\nabla\cdot(a(x)\nabla u(x))=f(x)\;\;x\in(0,1)^{2},$
$u(x)=0\;\;x\in\partial(0,1)^{2},$ (6)
where $u({\bf x})$ is the velocity of the flow, $a({\bf x})$ characterizes the conductivity of the media, and $f({\bf x})$ is the source function that can represent flow sources or sinks within the domain. In the experiment, our goal is to predict the solution $u$ given the external source $f$. To this end, we fixed the conductivity $a$, which is generated by first sampling a Gauss random field $\alpha$ in the domain and then applying a thresholding rule: $a({\bf x})=4$ if $\alpha({\bf x})<0$, otherwise $a({\bf x})=12$. We then used another Gauss random field to generate samples of $f$. We followed (Li et al., 2020b, ) to solve the PDE using a second-order finite difference solver and collected the source and solution at a $128\times 128$ grid.
A.2 Nonlinear Diffusion PDE
We next considered a nonlinear diffusion PDE,
$\partial_{t}u(x,t)$ $=10^{-2}\partial_{xx}u(x,t)+10^{-2}u^{2}(x,t)+f(x,t),$
$u(-1,t)$ $=u(1,t)=0,\;\;u(x,0)=0,$ (7)
where $(x,t)\in[-1,1]\times[0,1]$. Our objective is to predict the solution function $u$ given the source function $f$. We used the solver provided in (Lu et al.,, 2022), and discretized both the input and output functions at a $128\times 128$ grid. The source $f$ was sampled from a Gaussian process with an isotropic square exponential (SE) kernel for which the length scale was set to 0.2.
A.3 Eikonal Equation
Third, we employed the Eikonal equation, widely used in geometric optics and wave modeling. It describes given a wave source, the propagation of wavefront across the given media where the wave speed can vary at different locations. The equation is as follows,
$|\nabla u({\bf x})|=\frac{1}{f({\bf x})},\;\;{\bf x}\in[0,256]\times[0,256],$ (8)
where $u({\bf x})$ is the travel time of the wavefront from the source to location ${\bf x}$, ||$\cdot$|| ⋅ | denotes the Euclidean norm, and $f({\bf x})>0$ is the speed of the wave at ${\bf x}$.
In the experiment, we set the wave source at $(0,10)$. The goal is to predict the travel time $u$ given the heterogeneous wave speed $f$. We sampled an instance of $f$ using the expression:
$f({\bf x})=\max(g({\bf x}),0)+1.0,$
where $g(\cdot)$ is sampled from a Gaussian process using the isotropic SE kernel with length-scale $0.1$. We employed the eikonalfm library (https://github.com/kevinganster/eikonalfm/tree/master) that implements the Fast Marching method (Sethian,, 1999) to compute the solution $u$.
A.4 Poisson Equation
Fourth, we considered a 2D Poisson Equation,
$-\Delta u=f,\quad\text{in }\Omega=[0,1]^{2},\quad u|_{\partial D}=0.$ (9)
where $\Delta$ is the Laplace operator. The solution is designed to take the form, $u(x_{1},x_{2})=\frac{1}{\pi K^{2}}\sum_{i=1}^{K}\sum_{j=1}^{K}a_{ij}(i^{2}+j^{2})^{r}\sin(i\pi x_{1})\cos(j\pi x_{2})$, and $f(x_{1},x_{2})$ is correspondingly computed via the equation. To generate the dataset, we set $K=5$ and $r=0.5$, and independently sampled each element $a_{ij}$ from a uniform distribution on $[0,1]$.
A.5 Advection Equation
Fifth, we considered a wave advection equation,
$\frac{\partial u}{\partial t}+\frac{\partial u}{\partial x}=f,\quad x\in[0,1],\quad t\in[0,1].$ (10)
The solution is represented by a kernel regressor, $u({\bf x})=\sum_{j=1}^{M}w_{j}k({\bf x},{\bf z}_{j})$, and the source $f$ is computed via the equation. To collect instances of $(f,u)$, we used the square exponential (SE) kernel with length-scale $0.25$. We randomly sampled the locations ${\bf z}_{j}$ from the domain and the weights $w_{j}$ from a standard normal distribution.
A.6 Fatigue Modeling
Figure 4: Example of semi-elliptic surface crack on a plates (Merrell et al.,, 2024).
We considered predicting the SIF values along semi-elliptic surface cracks on plates, as shown in Fig 4. The SIF value can be viewed as a function of the angle $\phi\in[0,\pi]$, which decides the location of each point on the crack surface. The geometry parameters that characterize the crack shape and position were used as the input, including $a/c$, $a/t$ and $c/b$. In the operator learning framework, the input can be viewed as a function with three constant outputs. The dataset was produced via a high-fidelity FE model under Mode I tension (Merrell et al.,, 2024). Each data instance includes 128 samples of the SIF values drawn uniformly across the range of $\phi$.
A.7 Hyperparameter Settings
For the pseudo physics neural network $\phi$ — see (3) — we tuned the kernel size from {(3, 3), (5, 5), (7, 7), (9, 9)}. The stride was set to 1 and padding was set to “same” to ensure the output shape does not change. In the subsequent fully connected layers, we chose the number of layers from {3, 4, 5, 6}, and the layer width from {16, 32, 64}. We used GeLU activation. For FNO, we set the number of modes to 12 and channels to 32 (in the lifted space). We varied the number of Fourier layers from {2, 3, 4}. For DONet, in all the cases except Darcy Flow, the trunk net and branch net were constructed as fully connected layers. We varied the number of layers from {2, 3, 4} and the layer width was chosen from {30, 40, 50, 60}, with ReLU activation. For the case of Darcy flow, we found that DONet with only fully connected layers exhibited inferior performance. To address this, we introduced convolution layers into the branch net. We selected the number of convolution layers from {3,5,7}, and employed batch normalization and leaky ReLU after each convolution layer. To incorporate the learned pseudo physics representation into the training of FNO or DONet, we randomly sampled 200 input functions to construct the second loss term in (5). We set the maximum number of iterations to $10$ and selected the weight $\lambda$ from $[10^{-1},10^{2}]$. All the models were implemented by PyTorch (Paszke et al.,, 2019), and optimized with ADAM (Kingma and Ba,, 2014). The learning rate was selected from $\{10^{-4},5\times 10^{-4},10^{-3}\}$. The number of epochs for training or fine-tuning FNO, DONet and pseudo physics network $\phi$ was set to 500 to ensure convergence.
Table 3: Relative $L_{2}$ error in five operator learning benchmarks with richer data, where “PPI” is short for “Pseudo-Physics Informed”. The results were averaged from five runs.
(a) Darcy flow
Training size 600 1000
FNO 0.0093 $\pm$ 0.00015 0.0079 $\pm$ 0.00018
PPI-FNO 0.0087 $\pm$ 0.00040 0.0082 $\pm$ 0.00039
DONet 0.0540 $\pm$ 0.00064 0.0446 $\pm$ 0.00023
PPI-DONet 0.0415 $\pm$ 0.00077 0.0362 $\pm$ 0.00049
(b) Nonlinear diffusion
Training size 600 1000
FNO 0.0035 $\pm$ 0.0007 0.0028 $\pm$ 0.0004
PPI-FNO 0.0047 $\pm$ 0.00102 0.0042 $\pm$ 0.00096
DONet 0.0222 $\pm$ 0.00020 0.0187 $\pm$ 0.00023
PPI-DONet 0.0379$\pm$ 0.00088 0.0368 $\pm$ 0.00093
(c) Eikonal
Training size 600 1000
FNO 0.0193 $\pm$ 0.00017 0.0148 $\pm$ 0.00009
PPI-FNO 0.0199 $\pm$ 0.00016 0.0160 $\pm$ 0.00009
DONet 0.0460 $\pm$ 0.00021 0.0411 $\pm$ 0.00038
PPI-DONet 0.0475$\pm$ 0.00047 0.0436 $\pm$ 0.00034
(d) Poisson
Training size 600 1000
FNO 0.0045 $\pm$ 0.00006 0.0037 $\pm$ 0.00005
PPI-FNO 0.0046 $\pm$ 0.00005 0.0040 $\pm$ 0.00006
DONet 0.1786 $\pm$ 0.00398 0.1719 $\pm$ 0.00643
PPI-DONet 0.1409 $\pm$ 0.00292 0.1373 $\pm$ 0.00136
(e) Advection
Training size 600 1000
FNO 0.0943 $\pm$ 0.00177 0.0768 $\pm$ 0.00182
PPI-FNO 0.0819 $\pm$ 0.00092 0.0695 $\pm$ 0.00092
DONet 0.0913 $\pm$ 0.00074 0.0748 $\pm$ 0.00098
PPI-DONet 0.0732 $\pm$ 0.00143 0.0626 $\pm$ 0.00107
Table 4: Relative $L_{2}$ error in Poisson and Advection operator learning benchmarks, where “PPI” is short for “Pseudo-Physics Informed” and “PI” is truly “Physics Informed”. The results were averaged from five runs.
(a) Poisson
Training size 5 10 20 30
FNO 0.2340 $\pm$ 0.0083 0.1390 $\pm$ 0.0007 0.0895 $\pm$ 0.0008 0.0698 $\pm$ 0.0014
PPI-FNO 0.1437 $\pm$ 0.0062 0.0771 $\pm$ 0.0018 0.0544 $\pm$ 0.0009 0.0458 $\pm$ 0.0003
PI-FNO 0.1433 $\pm$ 0.0104 0.0718 $\pm$ 0.0015 0.0504 $\pm$ 0.0009 0.0429 $\pm$ 0.0004
(b) Advection
Training size 20 30 50 80
FNO 0.4872 $\pm$ 0.0097 0.4035 $\pm$ 0.0086 0.3019 $\pm$ 0.0085 0.2482 $\pm$ 0.0059
PPI-FNO 0.3693 $\pm$ 0.0099 0.3224 $\pm$ 0.0123 0.2236 $\pm$ 0.0075 0.1698 $\pm$ 0.0075
PI-FNO 0.3628 $\pm$ 0.0082 0.3205 $\pm$ 0.0121 0.$2$$\pm$ 0.0074 0.1668 $\pm$ 0.0057
Table 5: Relative $L_{2}$ error of using the learned back-box PDE network (3) to predict the input function $f$.
(a) Training size=10
Benchmark MLP Ours
Darcy Flow 0.1819$\pm$0.0026 0.1392$\pm$ 0.0080
Nonlinear Diffusion 0.0660$\pm$0.0069 0.0233$\pm$0.0005
Eikonal 0.0144$\pm$0.0009 0.0108 $\pm$ 0.0006
(b) Training size=30
Benchmark MLP Ours
Darcy Flow 0.1413$\pm$0.0013 0.0688$\pm$ 0.0032
Nonlinear Diffusion 0.0463$\pm$0.0022 0.0163$\pm$0.0002
Eikonal 0.0070$\pm$0.00005 0.0052 $\pm$ 0.0002
Table 6: The relative $L_{2}$ error with using different architectures of $\phi$ in pseudo-physics-informed (PPI) learning on Darcy flow benchmark.
(a) Predicting $f$ via $\phi$ with different architectures.
Training size 5 10 20 30
FNO 0.7229$\pm$0.0318 0.5759$\pm$ 0.0126 0.4257$\pm$ 0.0106 0.3160$\pm$ 0.0037
MLP 0.7169$\pm$0.0160 0.6598$\pm$ 0.0056 0.6464$\pm$ 0.0029 0.6277$\pm$ 0.0032
Ours 0.2285 $\pm$ 0.0147 0.1392 $\pm$ 0.0080 0.0898 $\pm$ 0.0046 0.0688 $\pm$ 0.0032
(b) Predicting $u$.
Training size 5 10 20 30
PPI-FNO with FNO as $\phi$ 0.5853$\pm$0.0153 0.3871$\pm$ 0.0124 0.2613$\pm$ 0.0190 0.1629$\pm$ 0.0064
PPI-FNO with MLP as $\phi$ 0.7262$\pm$0.0920 0.5516$\pm$ 0.0699 0.4568$\pm$ 0.0857 0.3983$\pm$ 0.1051
Standard FNO 0.4915 $\pm$ 0.0210 0.3870 $\pm$ 0.0118 0.2783 $\pm$ 0.0212 0.1645 $\pm$ 0.0071
Ours 0.1716 $\pm$ 0.0048 0.0956 $\pm$ 0.0084 0.0680 $\pm$ 0.0031 0.0642 $\pm$ 0.0010
Table 7: The relative $L_{2}$ error of PPI learning by incorporating different orders of derivatives. During the comparison with other operator learning methods, we used derivative orders up to 2 to run our method.
(a) Predicting $f$ via $\phi$.
Training size 5 10 20 30
order 0 0.7126$\pm$0.0131 0.5733$\pm$0.0208 0.4812$\pm$0.0399 0.3445$\pm$0.0182
order $\leq$ 1 0.2926$\pm$0.0118 0.2006$\pm$0.0047 0.1379$\pm$0.0051 0.1084$\pm$0.0053
order $\leq 2$ 0.2285$\pm$0.0147 0.1392$\pm$0.0080 0.0898$\pm$0.0046 0.0688$\pm$0.0032
order $\leq 3$ 0.2058$\pm$0.0192 0.1123$\pm$0.0039 0.0712$\pm$0.0021 0.0585$\pm$0.0030
(b) Predicting $u$.
Training size 5 10 20 30
order 0 0.6352$\pm$0.0673 0.4523$\pm$0.0621 0.3570$\pm$0.0658 0.2737$\pm$0.0643
order $\leq$ 1 0.3386$\pm$0.0259 0.2161$\pm$0.0083 0.1645$\pm$0.0114 0.1197$\pm$0.0132
order $\leq 2$ 0.1716$\pm$0.0048 0.0956$\pm$0.0084 0.0680$\pm$0.0031 0.0642$\pm$0.0010
order $\leq 3$ 0.2959$\pm$0.0381 0.1719$\pm$0.0213 0.1193$\pm$0.0158 0.0828$\pm$0.0054
Table 8: Parameter counts for FNO and DONet with PPI variations across different problems. The training size is 30.
Parameter count FNO PPI-FNO (increase) DONet PPI-DONet (increase)
Darcy-flow 1,188,353 1,229,476 (+3.46%) 2,084,704 2,125,827 (+1.97%)
Nonlinear-diffusion 1,188,353 1,197,220 (+0.75%) 824,501 833,368 (+1.08%)
Eikonal 1,188,353 1,197,220 (+0.75%) 824,501 833,368 (+1.08%)
Poisson 1,188,353 1,197,220 (+0.75%) 824,501 833,368 (+1.08%)
Advection 1,188,353 1,197,220 (+0.75%) 210,101 218,968 (+4.22%)
Appendix B Pointwise Error
For a further assessment, we conducted a fine-grained evaluation by visualizing the predictions and point-wise errors made by each method. In Fig. 7(a) and 7(b), we showcased the predictions and point-wise errors using PPI-DONet for Darcy Flow, PPI-FNO for nonlinear diffusion, respectively. Additional examples of predictions and point-wise errors are provided in Fig. 5, 8(a), 8(b), 9(a), and 9(b).
It is evident that without the assistance of the pseudo physics laws learned by our method, the ordinary DONet and FNO frequently missed crucial local structures, sometimes even learning entirely incorrect structures. For example, In Fig. 7(a) the first row, DONet missed one mode, while in the second and third row of Fig. 7(a), DONet failed to capture all the local modes. After incorporating the learned physics, DONet (now denoted as PPI-DONet; see the third column) successfully captures all the local modes, including their shapes and positions. Although not all the details are exactly recovered, the point-wise error is substantially reduced, particularly in those high error regions of the ordinary DONet; see the fourth column of Fig. 7(a). In another instance, as shown in Fig. 7(b), where the ordinary FNO (second column) captured the global shape of the solution, but the mis-specification of many local details led to large point-wise errors across many regions (fourth column). In contrast, PPI-FNO (third column) not only identified the structures within the solution but also successfully recovered the details. As a result, the point-wise error (fifth column) was close to zero everywhere. Additional instances can be found in Fig. 8(a), the first three rows illustrate that ordinary FNO (trained with 5, 10, and 20 examples, respectively) estimates an entirely incorrect structure of the solution, indicating that the training data is insufficient for FNO to capture even the basic structure of the solution. In contrast, after fine-tuning with our learned physics laws from the same sparse data, PPI-FNO accurately figured out the solution structures and yielded a substantial reduction in point-wise error across nearly everywhere. The point-wise error became uniformly close to zero. With 30 examples, the ordinary FNO was then able to capture the global structure of the solution, but the details in the bottom left, bottom right, and top right corners were incorrectly predicted. In comparison, PPI-FNO further recovered these details accurately.
Collectively, these results demonstrate that the pseudo physics extracted by our method not only substantially boosts the overall prediction accuracy but also better recovers the local structures and details of the solution.
Figure 5: Examples of SIF prediction of FNO and PPI-FNO trained with 600 examples.
(a) PPI-FNO: learning
(b) PPI-DONet: learning
(c) PPI-FNO: $\lambda$
(d) PPI-DONet: $\lambda$
Figure 6: Learning curve of PPI-FNO (a) and PPI-DONet (b) on Eikonal with 30 training examples. Shown in (c) and (d) is how the weight $\lambda$ of “pseudo physics” affects the operator learning performance. The horizontal lines in (c) and (d) are the relative $L_{2}$ error of standard FNO and DONet, respectively.
(a) PPI-DONet: Darcy Flow
(b) PPI-FNO: nonlinear diffusion
Figure 7: Examples of the prediction and point-wise error of PPI-DONet and PPI-FNO on Darcy Flow and nonlinear diffusion, respectively. From top to bottom, the models were trained with 5, 10, 20, 30 examples.
(a) PPI-FNO: Eikonal
(b) PPI-DONet: Eikonal
Figure 8: Examples of the prediction and point-wise error of PPI-FNO and PPI-DONet on Eikonal. From top to bottom, the models were trained with 5, 10, 20, 30 examples.
(a) PPI-FNO: Darcy Flow
(b) PPI-DONet: Nonlinear Diffusion
Figure 9: Examples of the prediction and point-wise error of PPI-FNO and PPI-DONet on Darcy Flow and Nonlinear diffusion, respectively. From top to bottom, the models were trained with 5, 10, 20, 30 examples.
Appendix C Limitation and Discussion
Our current method cannot learn PDE representations for which the input function $f$ is the initial condition. In such cases, the mapping from the solution function to the initial condition requires a reversed integration over time, hence we cannot decouple the derivatives. To address this problem, we plan to explicitly model the temporal dependencies in the PDE representation, such as via the neural ODE design (Chen et al.,, 2018).